# Effective Medium Theory for Time-modulated Subwavelength Resonators

### v1

@@ 6 Concluding Remarks
We then proceeded with the case of the time-dependent material parameter $\kappa$. In this paper, we assumed the parameter $\rho$ to be static, since in [1] we proved that the resulting wave field does not depend on $\rho$ at leading order. As in the static case, we first derived the scattering matrix. However, for a time-modulated material parameter, this is a $2(2K+1)\times 2(2K+1)$ matrix. This is a direct consequence of the mode coupling that arises from the modulation in time. Analogously to the static case we then derived a point interaction approximation, which ultimately led to a characterisation of the effective field through a system of coupled Lippmann-Schwinger equations (4.35). These integral equations furnished the homogenised equations given by (4.36). In contrast to the static case, the homogenised equations modelling a time-modulated metamaterial are a system of coupled integro-differential equations.
A crucial assumption to obtain a time-modulated effective medium theory is that the incident frequency $\omega$ is slightly away from the resonant frequency of the components. For a frequency at resonance, the scattering coefficients $g$ are of order one. Even for a large number $N$ of resonators, individual resonator gives a strong contribution to the total field, and the limit does not exist.
@@
We consider the results proven in this paper to be the basis for many new breakthroughs in the exploration of time-modulated metamaterials. Since our results can be easily generalised to higher dimensions, the results herein proved are also of great significance to two- and three-dimensional spacetime metamaterials. Moreover, we showed that time-dependent metamaterials at a resonant frequency cannot be treated as an effective medium.

### v3

@@ 6 Concluding Remarks
We then proceeded with the case of the time-dependent material parameter $\kappa$. In this paper, we assumed the parameter $\rho$ to be static, since in [2] we proved that the resulting wave field does not depend on $\rho$ at leading order. As in the static case, we first derived the scattering matrix. However, for a time-modulated material parameter, this is a $2(2K+1)\times 2(2K+1)$ matrix. This is a direct consequence of the mode coupling that arises from the modulation in time. Analogously to the static case we then derived a point interaction approximation, which ultimately led to a characterisation of the effective field through a system of coupled Lippmann-Schwinger equations (4.38). These integral equations furnished the homogenised equations given by (4.39). In contrast to the static case, the homogenised equations modelling a time-modulated metamaterial are a system of coupled differential equations.
A crucial assumption to obtain a time-modulated effective medium theory is that the incident frequency $\omega$ is slightly away from the resonant frequency of the components. For a frequency at resonance, the scattering coefficients $g$ are of order one. Even for a large number $N$ of resonators, each individual resonator makes a strong contribution to the total field, and the limit does not exist. In summary, we showed that time-dependent metamaterials at a resonant frequency cannot be treated as an effective medium.
@@
We consider the results proven in this paper to be the basis for many new breakthroughs in the exploration of time-modulated metamaterials, similarly to achievements regarding static metamaterials [7]. Since we expect our results to generalise to higher dimensions (see Remark 4.11), the results herein proved are also of great significance to two- and three-dimensional spacetime metamaterials.